OpenStudy (anonymous):

If a number is subtracted from it's square the result is 2. Find the number.

Parth (parthkohli):

\( \color{Black}{\Rightarrow x^2 - x = 2}\)

Parth (parthkohli):

\( \color{Black}{\Rightarrow x^2 - x - 2 = 0}\) \( \color{Black}{\Rightarrow x^2 + x - 2x - 2 =0}\) \( \color{Black}{\Rightarrow x(x + 1) - 2(x + 1) = 0}\) \( \color{Black}{\Rightarrow ( x - 2)(x + 1) = 0}\) \( \color{Black}{\Rightarrow x = 2,-1 }\)

OpenStudy (anonymous):

So it's 2 or -1?

Parth (parthkohli):

Yes.

OpenStudy (anonymous):

The square of -1 is -1 and -1 minus -1 is 0. Not 2.

Parth (parthkohli):

\( \color{Black}{\Rightarrow (-1)^2 - (-1)}\) \( \color{Black}{\Rightarrow (-1)^2 + 1}\) \( \color{Black}{\Rightarrow (-1)(-1) + 1}\) \( \color{Black}{\Rightarrow 1 + 1}\) \( \color{Black}{\Rightarrow 2}\)

OpenStudy (anonymous):

i guess